OpenStudy (anonymous):

Write y = x + 5 in standard form using integers. A. –x – 6y = 30 B. 6x – y = 30 C. –x + 6y = 30 D. –x + 6y = 5

OpenStudy (anonymous):

I am stuck

OpenStudy (anonymous):

I need help with this one

OpenStudy (anonymous):

Subtract x on both sides -x+y=5 But your y-int should be 5 so multiply y by 6 and 5 by 6 To get -x+6y=30 So C.

OpenStudy (anonymous):

Thanks so much!! I was so stuck..

OpenStudy (anonymous):

Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers.

OpenStudy (anonymous):

so is it c?

OpenStudy (anonymous):

Well our A is technically "-1x" Wait for some_someone finishes maybe they know what to do...

OpenStudy (anonymous):

\[ –x – 6y = 30 \] \[y = -\frac{ x }{ 6 }-5\] ------------------------- \[6x – y = 30 \] \[y = 6 x-30\] \[y = x-5\] -------------------------- \[–x + 6y = 30 \] \[y = \frac{ x }{ 6 }+5\] -------------------------- \[–x + 6y = 5\] \[y = \frac{ x }{ 6 }+\frac{ 5 }{ 6 }\]

OpenStudy (anonymous):

So, I would say -- B. 6x – y = 30

OpenStudy (anonymous):

I have another question

OpenStudy (anonymous):

Match the equation with its graph. –4x – 2y = 8

OpenStudy (anonymous):

Slope-int. form is y=-2x-4 Because –4x – 2y = 8 Add 4x on both sides -2y=4x+8 Divide by -2 y=-2x-4

OpenStudy (anonymous):

you rock @RoseDryer

OpenStudy (anonymous):

For your first question though. The answer should be x - y = -5 Which isn't one of the choices...

OpenStudy (anonymous):

I have another question Does the equation represent a direct variation? If so, find the constant of variation. 5x = -3y

OpenStudy (anonymous):

I suck at direct variation.

OpenStudy (anonymous):

lol I suck at math